Calculate the Kc of the reaction below if the equilibrium concentrations of Ag and IO3 are both 1.78×10-⁴M

1
Expert's answer
2022-07-04T01:01:04-0400


AgIO3 <=> Ag+ + IO3-


1 - 1.78*10-4 1.78*10-4 1.78*10-4



"K_c = ~\dfrac {[Ag^+][IO_3^-]}{[AgIO_3]}=~\dfrac {(1.78*10^{-4})^2}{1-1.78*10^{-4}}=~ 3.17*~10^{-8}"



Answer: 3.17*10-8